F=(9/5)*C (degrees)+32
Find the F temp if it is 32degrees Celsius
ALso what if the temp is -40degreesF, what woulld it be in celsius

Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16949)

F=(9/5)*C (degrees)+32
plug the value of C in the equation
F=(9/5)*32+32
F=57.6+32
F=89.6 deg
......
ALso what if the temp is -40degreesF, what woulld it be in celsius
-40=(9/5)*C +32
add -32
-40-32= 9/5*c
-72=9/5*c
5*-72/9 =C
c= -40 deg
...
m.ananth@hotmail.ca

Answer by ewatrrr(24785)


Hi,
F=(9/5)*C (degrees)+32
5/9(F° - 32°)= C
Find the F temp if it is 32° Celsius
5/9(32° - 32°)= C
0° = degrees celsius
substituting -41° for F will give you the degrees celsius for that.
5/9(F° - 32°)= C